You have 75 mL of a solution with [Zn2+] = 0.031 M. The pH of this solution = 8. You titrate this solution with 0.055M EDTA. Remember that pMn+ = -log(Mn+)
a) What is true at the equivalence volume for any titration?
b) Calculate the Ve (the equivalence volume) for this titration?
c) Calculate the pZn2+ after addition of 20 mL EDTA
d) Calculate the pZn2+ at the equivalence volume
e) Calculate the pZn2+ after addition of 50 mL EDTA
1
Expert's answer
2020-04-09T13:09:49-0400

a) At the equivalence point of any titration the number of equivalents (in moles) of the titrant is equival to the number of equivalents (in moles) of the titrated substance.


b) The titration is based on complex formation:

Zn2+ + EDTA4- = ZnEDTA2-

From the definition of equivalence point, [EDTA4-]*Ve = VZn*[Zn2+]

Ve = 75*0.031/0.055 = 42.27 mL


c) after the reaction with 20 mL of EDTA solution, the rest amount of Zn2+ is

0.075*0.031 - 0.020*0.055 = 0.001225 mole,

the final volume of solution is 0.075 + 0.020 = 0.095 L,

concentration of [Zn2+] = 0.001225/0.095 = 0.013 M,

pZn2+ = 1.89


d) at the equivalence point the concentration of Zn2+ is only due to the ZnEDTA2- dissociation equilibrium, no excess concentration of Zn2+ and EDTA2- are in the solution. The logarithm of formation constant of the complex is lgKf = 16.5.

Kf = [ZnEDTA2-]/([Zn2+]*[EDTA4-])

Concentration of [ZnEDTA2-] equals to that of initial [Zn2+] concentration (or the added [EDTA4-] concentration) with an account of dilution:

[ZnEDTA2-] = 0.031*75/(75+42.27) = 0.020 M

Equilibrium concentrations of [Zn2+] and [EDTA4-] are equal, hence:

[Zn2+] = sqrt([ZnEDTA2-]/Kf) = sqrt(0.020/1016.5) = 10-9 M,

pZn2+ = 9


e) after addition of 50 mL EDTA (above the equivalence volume) the concentration of Zn2+ is due to complex ZnEDTA2- dissociation in presence of excess of EDTA4-. The amoun of free EDTA4- ligand is 0.050*0.055 - 0.04227*0.055 = 0.000425 mole, the final volume is 0.075 + 0.050 = 0.125 L, the concentration of excess [EDTA4-] = 0.000425/0.125 = 0.0034 M.

The concentration of [ZnEDTA2-] equals to that that of initial [Zn2+] with an account of dilution:

[ZnEDTA2-] = 0.031*75/(75+50) = 0.0186 M.

From the expression for formation constant we find that:

[Zn2+] = [ZnEDTA2-]/(Kf*[EDTA4-]) = 0.0186/(1016.5*0.0034) = 10-15.76

pZn2+ = 15.76